A block-spring system undergoes simple harmonic motion with an amplitude A. 3.1 If the mass is doubled but the amplitude remains unchanged, how will this affect the total energy of the system? 3.2 Can the displacement and the acceleration of the mass be In the same direction? Explain.
Answer #1

1) Total energy remains the same.

Total energy is given by:

U = 1/2*k*x_max^2 = 1/2*k*A^2

So, its independent of mass.

2) No. The acceleration is proportional to the displacement but is directed in the opposite direction.